9. The reaction 2A → B is second order with a rate constant of 51.0/M·min at 24°C.

(a) Starting with [A]0 = 9.50 × 10−3M, how long will it take for [A]t = 3.10 × 10−3M?

______ min

(b) Calculate the half-life of the reaction.

_______ min

10. The thermal decomposition of phosphine (PH3) into phosphorus and molecular hydrogen is a first-order reaction: 4PH3(g) → P4(g) + 6H2(g) The half-life of the reaction is 35.0 s at 680°C.

a) Calculate the first-order rate constant for the reaction:

_______ s−1

b) Calculate the time required for 78.0 percent of the phosphine to decompose:

________ s
Answer #1

2A B second order reaction rate constant = K = 5/01 Momin K= 51.0 Mlmin ♡ Chlo = 9.50 X10-3M [A]+= 3,10 8 10-3M time requiredkindly post different questions separately. please rate